You are going to paint a wall in your room. The wall is 12 feet long and 8 feet high. A window in the wall is 3 feet wide and 4 feet high. Find the area that you will paint.

A rectangle has a length of 4 inches and width of 3 inches. How would the area change if you doubled both dimensions? Explain your reasoning.

Answer by checkley79(3341)

12*8-3*4
96-12=84 ft^2 ans.
------------------------
4*3=12 in^2 is the original area.
(2*4)(2*3)
8*6=48 in^2 is the expanded area.
The area is 4 times the original area because 2*2=4